### v1

# Event-Based Simulation of Stochastic Memristive Devices for Neuromorphic Computing

Abstract
In this paper, we build a general model of memristors suitable for the simulation of event-based systems, such as hardware spiking neural networks, and more generally, neuromorphic computing systems. We extend an existing general model of memristors - the Generalised Metastable Switch Model [1] - to an event-driven setting, eliminating errors associated discrete time approximation, as well as offering potential improvements in terms of computational efficiency for simulation. We introduce the notion of a volatility state variable, to allow for the modelling of memory-dependent and dynamic switching behaviour, succinctly capturing and unifying a variety of volatile phenomena present in memristive devices, including state relaxation, structural disruption, Joule heating, and drift acceleration phenomena. We supply a drift dataset for titanium dioxide memristors and introduce a linear conductance model to simulate the drift characteristics, motivated by a proposed physical model of filament growth. We then demonstrate an approach for fitting the parameters of the event-based model to the drift model.
Index Terms:
memristors, neuromorphic computing, event-based models, volatility
I Introduction
The modelling of memristive devices is an important component of the development of novel architectures and systems based on these devices, including neuromorphic and event-based computing systems. Models can be classified along several dimensions: their generality, whether they are data-driven or physics based (or both), the inclusion of stochastic or non-ideal behaviour, and suitability for simulation in terms of their computational efficiency (compactness) in the relevant simulation context. The last of these dimensions can be thought of as key - since the goal of a model is to enable the prediction of behaviour so as to allow for the design of strategies to utilise the behaviour of the devices to our advantage. The specific goal of our modelling approach is the development of event-based neuromorphic computing simulations. Accordingly, this paper aims to extend existing modelling approaches for memristive devices to an event-based setting to allow the models to be used for neuromorphic computing simulations, by introducing general modelling principles, applicable to a broad range of memristive devices.
Figure 1: A model’s use is governed by its predictive ability in a particular setting. We focus on memristor modelling appropriate for use in a neuromorphic computing setting, introducing event-based simulation and the use of a volatility state variable to enable the model to be used to simulate key behaviours in this setting.
I-A Memristors
The term memristor has come to encompass a great many types of device, since its proposition as the fourth fundamental passive circuit element by Chua [2]. The original formulation proposed an equation linking the charge to the flux through a (polynomial) relationship, which resulted in an instantaneous resistance determined by the ratio of the voltage to the current, and dependent (in some way) on the instantaneous charge present in the device. The rate of change of the device’s resistive state was purely a function of the state itself and the magnitude of the voltage applied across the device (or the current flowing into it).
To expand the scope of memristors, Chua & Kang published a paper introducing the notion of memristive systems [3]. Memristive systems were defined purely mathematically, with no restrictions on the nature of the physical quantities, but only the relationship between them, also allowing for one or more state variables to exist in the device. Despite its generality, this model was unable to capture autonomous state evolution dynamics (later addressed by other models [4]), and was far removed from the physical world, by design, to allow it to be as general as possible, limiting its use for modelling real devices.
A change in the definition of the term “memristor” was later ignited by the identification of reasonably reliable resistive switching behaviour in a titanium dioxide thin film device [5], and subsequently fuelled by the broadening of the definition by Chua [6]. The types of device that have been described as memristors, commonly grouped according to the switching mechanism, include Phase Change Memory (PCM), Electrochemical Metallization (ECM), Valence Change Mechanism (VCM), Thermochemical Mechanism (TCM), Electronic (including Mott insulator and electron trapping devices), magnetoresistive memories (i.e. Spin-Transfer Torque (STT) or spintronic memristors), biological memristors, Micro Electromechanical Systems (MEMS). The switching mechanism is often correlated with the material type. For example, VCM and TCM switching devices are based on the movement of oxygen vacancies through the active layer of the material, while PCM devices are often constructed from a chalcogenide glass such as Ge-Sb-Te (GST).
I-B Event Based Systems
Event-based systems are systems where changes to the state are driven by the time of events occurring, rather than according to a pre-defined sample time. Such systems can efficiently process information that is sparse in time. Systems that process this kind of information would need to be simulated using finite difference equations, which may be both inefficient due to the need to simulate each and every time-step, and also inaccurate, with inaccuracies that compound over time due to the use of a non-infinitesimal (finite) time-step for simulation.
Our motivation for introducing an event-based simulation framework is to allow for the simulation of memristors in event-based settings, most notably, spiking neural networks implemented on hardware, in the context of the growing research area of neuromorphic computing (computing that aims to use principles from neuroscience to improve the processing of information) [7]. Indeed, the notion of event-based processing is inspired by the function of the human brain. Mmristors have been touted as potential neuromorphic computing elements, due to their ability to process information in-place, circumventing the Von Neumann bottleneck. This bottleneck refers to the separation of storage and processing and the associated high energy and computation time costs of transferring information from the ‘store’ (storage) to the ’mill’ (processor) and back during computation [8]. Circumventing the Von Neumann bottleneck offers the potential for massively parallel processing, akin to that possible with biological neurons.
It is clear that in order to implement event-based computation in hardware with memristors, it would be more natural to use models which are suited to event-based simulation settings.
I-C Volatility
One key aspect of information processing in the human brain is the split of memory into long-term and short-term, often labelled as long term potentiation (LTP) and short term potentation (STP). The state variable in the memristor model can capture long-term memory changes, while the possibility for autonomous state evolution captures short-term memory change to some degree. In order to enrichen the model’s ability to describe STP phenomena, we account for the possibility of short term memory outside of the state variable through the inclusion of a volatility state variable. This variable allows for the capture of a variety of physical phenomena in various memristive systems, including Joule heating in thin film metal oxide devices, non-linear drift in metal oxide devices, and state relaxation in phase change memories.
Several different phenomena in both metal oxide devices and phase change memories rely on the principle of autonomous state evolution. In metal oxide devices, the autonomous degradation of the conductive filaments results in the phenomenon of resistive drift. In phase change memories, the phenomenon of structural relaxation results in a change of the device’s resistance over time. As a result, we also extend the modelling framework by introducing a volatility state variable, as a unified way of describing the autonomous state evolution behaviour seen in a variety of memristive systems.
I-D Contributions
We introduce two major contributions to memristive modelling. Our first contribution is to look at a new simulation paradigm for event-based systems which is natively event-driven, based on the evaluation of time-to-fire for multiple parallel Poisson point processes, while also allowing for continuous-time changes - such as changes in temperature or bias - to be incorporated. We extend an existing generic memristor model, the Generalized Metastable Switch Model (GMSM) [1, 9] to the event-based setting, in order to address problems associated with the time-stepped simulation, where error is inevitable and proportional to the size of the timestep. We modify the rate calculations of the model, to base them on a more physically generic Boltzmann probability density function. We also generalise the model through the separation of the resistance and the state variable via a readout equation.
Our second modelling contribution is to generalise a variety of nonlinear, temporary and autonomous evolution phenomena observed in different types of memristive device - thin film Metal-Insulator-Metal devices and phase change memories - and propose the use of a single variable called the “volatility state variable” to singularly describe all these phenomena.
In addition, as part of our modelling approach, we propose the separation of resistance, or other output variable, calculations using a readout equation. This approach separates the underlying memristive state variable from the resistance, through the resistance readout equation, allowing for simplification of the model for the switching behaviour.
Following the development of the model, we provide a real resistive drift dataset for a titanium oxide memristive device and propose a linear conductance model to model the resistive drift in this type of device, based on a filament growth model. We then demonstrate the ability of the model to replicate the resistive drift phenomenon present under zero-bias.
Notation We use parameters $E_{x}$ to denote an energy barrier height in terms of an electron, where $V_{x}=E_{x}/q$ is the corresponding equivalent potential difference. We use the notation $\mathcal{X}[i]$ to denote the $i$’th element in the ordered set $\mathcal{X}$.
II Event-Based Metastable Switching Model
In this section, we introduce the proposed modelling framework. We first propose an extension to the Generalised Metastable Switch Model [1] - a general model for memristive devices - and examine a modified metastable switch model based on continuous time probabilistic point processes, rather than discrete time Bernoulli random variables, in order to make the model event based. We then incorporate the so-called “volatility state variable” to allow for richer modelling of physical phenomena such as autonomous state evolution, Joule heating, and nonlinear drift, which in turn allows for the characterisation of certain short term memory behaviours in memristive devices.
II-A Limitations of Discrete-Time Stochastic Switching Models
Many models for memristive devices have been proposed; however, only a handful of these take into account the stochasticity inherent in the devices [10, 11], with the majority making deterministic assumptions.
One general memristor model that includes stochasticity as a fundamental feature of the model, is Molter & Nugent’s “Generalized Metastable Switch Model” (GMSM) [1]. It models the memristor as a number of parallel metastable switches, each allowed to take one of two different states - the “high resistance” or “low resistance” state (see Figure 2), with stochastic transitions occurring every unit time with probabilities that are dependent on the temperature and bias.
Figure 2: The metastable switch model. The conductance is a linear function of the number of switches (parallel conducting channels) that are in the “on” state at any one time.
The GMSM bases its switching mechanism upon a Fermi-Dirac-like activation function (a logistic function), based on the temperature and the bias, which dictates the rate at which switching events occur. Two switching processes are modelled: low resistance to high resistance, and high resistance to low resistance. The probability of a single metastable switch transitioning from the low resistance state to the high resistance state in a given timestep $\Delta T$ is denoted by $P_{A}$, while the probability of a transition from the high resistance state to a low resistance state in $\Delta T$ is denoted by $P_{B}$ These two probabilities are given by
$P_{A}$ $=\alpha\frac{1}{1+e^{\beta(V-V_{A})}},$ (1)
$P_{B}$ $=\alpha\frac{1}{1+e^{\beta(-V+V_{B})}},$ (2)
where $\beta\triangleq\frac{q}{k_{B}T}$ is a physical parameter that introduces temperature-dependence and is the inverse of the thermal voltage ($k_{B}$ is Boltzmann’s constant and $q$ is the electronic charge), $\alpha=\frac{\Delta t}{t_{c}}$ is a ratio of the timestep to the characteristic timescale $t_{c}$ a parameter of the device.
The GMSM models the device’s state evolution over time by considering the number of switches in either the high or low state and using the respective transition probabilities to define two binomial distributions that can be used to simulate the number of respective transitions of switches from low resistance to high, or vice versa for the second distribution, over a time-step. The normal approximation to a binomial random variable can be used, as $N$ (the number of metastable switches) becomes large, to develop a continuous stochastic model.
We note that the approximation in Equation 1 is valid only for small values of $\Delta t$. If we let the value of $\Delta t$ grow large, then errors associated with discretising the timestep begin to emerge. The larger the value of $\Delta t$, the more significant a change will have occurred in the parameters used to calculate the transition probabilities, leading to modelling inaccuracies. This is a problem for any discrete-time approximation of a continuous-time problem i.e. any finite difference method for approximating a continuous-time differential equation. Error-free modelling can only be achieved in the limit as the timestep vanishes, i.e., $\Delta t\rightarrow 0$.
In order to address the discrete-time errors associated with the GMSM, as well as being motivated by the use of memristors in event-based neuromorphic computing systems, we consider modelling the state transitions not in terms of a transition probability per timestep, but rather through a (Poisson) point process, whose rate parameter dictates the frequency of switching events. This approach switches from a timestep-based evaluation to an event-based evaluation, where we model the changes to the device’s state as and when they occur. This removes the discretisation error. It also offers potential benefits from a computational efficiency perspective, since the frequency of model updates are tied to frequency of change of the memristive state. Next, we specify the model dynamics and modelling principles.
II-B State Modelling
In keeping with the terminology of the GMSM, we use the terms “switches”, or “metastable switches”, to refer to the state variables of the model (the volatility is separately dealt with). For the purposes of this work, we assume that both the number of state variables ($N$) and the multiplicity of the individual state spaces are finite. Specifically, we consider the case where $k=2$ i.e. binary switching. The extension to the continuous setting is left as future work. We use $n$ to denote the number of switches in the state that is conducive to conduction, which we call the “up state”, the “high conductance state”, or the “low resistance state”. Then, $m=N-n$ denotes the number of switches in the state that is not conducive to conduction, which we call the “down state”, the “low conductance state”, or the “high resistance state”. Switching from a lower to a higher resistance state is termed “up” switching, while switching in the opposite direction is termed “down” switching, reflecting the respective increase or decrease in the resistance. Note that due to the deterministic relationship between $m$ and $n$, there is effectively only one state variable, yet we make use of both for notational convenience.
The physical mechanism motivating the use of $N$ parallel, binary, metastable switches used as a model for memristive devices could be one of the following: the number of oxygen vacancies in the bulk of the material vs. those which are attached to metal ions, where reduction has not occurred; the number of vacancies/ions that are present in the (main) conductive filament, versus those which are not, as per the popular, single dominant conductive filament model; or the number of parallel conducting pathways running through the device, vs. those that are not formed.
We detail the inclusion of an additional state variable that models the volatility in a device, motivated by physical evidence for similar phenomena in multiple resistive switching devices, in Section II-E1. This additional variable does not affect the readout equation, but may affect the switching rates. The motivation for having a single state variable that impacts the readout equation is that this is in keeping with the spirit of most resistive switching memories, whose state is thought to be closely tied to their instantaneous resistance. In addition, it has been shown that a single state variable is sufficient to explain the behaviour of various memristive devices [5, 12].
II-C Readout Equation
Note that the voltage-current relationship for a memristive device may not always be instantaneously linear. Indeed, an exponential or $\sinh$ relationship is possible depending on the nature of the device being considered [1]. Different behaviour can be considered by relating the memristive state variable, $n$, to a separate quantity such as the parameter for a Shockley diode equation, or the scaling factor for a $\sinh$ voltage-current relation.
Rather than requiring that the resistance be linearly linked to the state variable, we instead use a readout equation to provide flexibility in simulation. The readout equation takes as input the memristive state variable $n$, and outputs a resistance. This allows for a consistent description of the state, and allows for the separation of measurement errors from the state evolution dynamics. In Section 10, we see an example of a thresholding, non-linear readout equation, that links the state variable to the voltage-current characteristics in a non-linear way.
II-D Switching Mechanism
The GMSM models the metastable switching dynamics through a set of $N$ parallel Bernoulli random variables (one for each switch) for each timestep. As noted, our modelling approach differs in that it models the switches as independent Poisson point processes happening in parallel. The $N$ independent Bernoulli random variables are then combined into two binomial random variables - or, in the continuous approximation, a normal distribution - that measure the state change for all switches in a single timestep $\Delta T$.
We instead approach the modelling by replacing the Bernoulli probabilities with rates for the independent Poisson point processes. We use the convention that the subscript up refers to parameters describing the up transition (and the same, respectively, for subscript down and down transitions).
The rate parameter for the Poisson process of a single metastable switch is referred to as the base rate and denoted by either $\lambda^{\prime}_{\text{up}}$ or $\lambda^{\prime}_{\text{down}}$. In each case, its value is determined by a Boltzmann activation function (see Section II-D1 below) that is dependent on the voltage bias and the temperature, as well as static parameters. To obtain the overall rate (the cumulative rate of all up or down switching events), we simply scale the base rate by the number of switches in the origin state - $m$ for $\lambda^{\prime}_{\text{up}}$, or $n$ for $\lambda^{\prime}_{\text{down}}$, giving us the cumulative up and down switching parameters, $\lambda_{\text{up}}$ and $\lambda_{\text{down}}$. We use the convention that a positive voltage makes it more probable for the device to switch to a higher resistive state, while a negative voltage makes it more probable for the device to switch to a lower resistive state.
Equation 3 below, gives the Poisson rate of a single switch transitioning from the “high” - or “open circuit” - state, $m$, to the “low” - or “conducting” - state, $n$: $\lambda^{\prime}_{\text{down}}$. Note that in the equations, two physical constants are also used: Boltzmann’s constant, $k_{B}$, and the electronic charge, $q$. As a result, Equation 4 gives the rate parameter for the overall up switching process.
$\lambda^{\prime}_{\text{up}}(t)$ $=\exp{\left(-\frac{E_{a}-kV(t)q/2-E_{\text{off}}/2}{k_{B}T(t)(1+\rho(t))}\right)},$ (3)
$\lambda_{\text{up}}(t)$ $=n(t)\lambda^{\prime}_{\text{up}}(t),$ (4)
where the parameters are as follows:
  1. 1.
    $E_{a}=qV_{a}$ is the activation energy (the energy barrier height). This increases or decreases both the up and down switching rates.
  2. 2.
    $E_{\text{off}}=qV_{\text{off}}$ (added, rather than subtracted, for $\lambda^{\prime}_{down}$) is the energy offset that leads to a shift of the equilibrium point from equal numbers of states being in the high or low resistive state. This adjusts the balance of switching rates under zero bias.
  3. 3.
    $k$ is a constant factor used to convert an applied bias to the effective voltage used for determination of the effect voltage and subsequent shift in the barrier height, accounting for material-specific response factors, such as lattice periodicity.
  4. 4.
    $V(t)$ is the voltage across the device at time $t$, which will be added for $\lambda^{\prime}_{down}$ as opposed to being subtracted. This gives an additional energy offset when scaled by $q$.
  5. 5.
    $T(t)$ is the time-dependent temperature variable in degrees Kelvin. This may be influenced by the history of applied voltages and the device state history, which directly impact the rate of switching, as well as the equilibrium point when $E_{\text{off}}$ is not 0.
  6. 6.
    $\rho(t)$ has a similar impact to the temperature and is the volatility state variable, that may also be dependent on the time and voltage/state history.
In this model, application of a voltage bias, $V(t)$, lowers or raises the activation energy barrier in one direction, simultaneously raising or lowering it in the other direction. This is illustrated in Figure 3. Note also that, for the purposes of simplicity, in this paper, we assume, without loss of generality, that $k=1$.
Figure 3: Illustration of the effect of the different parameters on the Boltzmann barrier heights in the forward and reverse switching directions. In a) we see the standard barrier, unmodulated by an offset or an applied voltage. b) shows the barrier modified by an energy offset $E_{\text{off}}$. c) shows the energy barrier modulated by the offset in addition to the modulation of the barrier resulting from the application of a voltage bias, $E_{v}$. The relative energy barrier height for each local minimum determines the respective Poisson switching rates $\lambda^{\prime}_{\text{up}}$ and $\lambda^{\prime}_{\text{down}}$.
The volatility state variable is an additional state variable that is introduced to incorporate stateful effects on the evolution of the resistive state, discussed further in Section II-E.
II-D1 Boltzmann Activation Function
To determine the rates of the Poisson processes, we made use of a Boltzmann occupation function, rather than the Fermi-Dirac occupation function (logistic function) used to determine the switching probabilities in the GMSM. We make use of the Boltzmann distribution due to it being a good model of probabilistic behaviour in microscopic thermodynamic systems, modulated by energy barriers.
In the case of a metal oxide memristor, we envision ions hopping over energy barriers, mediated by voltage bias and temperature conditions, resulting in the movement of ions into more or less favourable locations for conduction (modulating the instantaneous resistance). The nature of the ions is not crucial: for example, they could be metal ions (as in ECM devices), or oxygen vacancies (as in VCM devices). There are many possible mechanisms that could drive the movement of ions in an ionic movement memristive device. Filamentary, interfacial, and bulk switching (e.g. Mott insulator-metal transition through ionic doping/trapping) mechanisms have been proposed. Such a physical switching mechanism motivates the use of a Boltzmann modelling approach. The use of sigmoid-like or Boltzmann-like activations is common in models of memristor state evolution, and a sigmoid activation is used in the GMSM simply to bound the probability of state transitions between 0 and 1 [1].
$\sinh$ switching behaviour (for an applied voltage) naturally arises out of modelling that uses an energy-barrier assumption with a height and hopping rate that is exponential in the applied voltage. $\sinh$-switching behaviour is something that has been observed in a variety of memristive devices and attributed to phenomena such as exponential ionic drift for high fields, or Joule heating effects [12, 13]. This can be observed by subtracting the instantaneous rate of upward transitions (transitions from the down state to the up state) from the downward transitions (transitions from the up state to the down state) as follows.
In [13], Strukov & Williams note that for the rigid point ion model, the energy barriers in the forward and reverse directions for the drift of ions in a lattice would be distorted, through the application of an electric field, by ±$\pm q\mathcal{E}a/2$, where $\mathcal{E}$ is the field, and $a$ is the periodicity of the lattice. The characteristic length $a$ need not be lattice periodicity, therefore we simply combine the impact of the lattice periodicity, specific to the rigid point ion model, and the electric field into a parameter $V=\mathcal{E}a$, thus the distortion will be ±$\pm qV/2$, respectively.
The average energy of any ion will be approximately the thermal voltage $V_{T}=k_{B}T$, thus providing a mechanism for thermally accelerated drift, through Joule heating effects.
Based on this, if we assume that the rate of hopping upwards or downwards for each “switch” is given by a rate equation based on the Boltzmann probability density [^1] The Boltzmann probability density is a common model for the probability density of particles at thermal equilibrium., then we can say that the rate of transition from the high resistance state to the low resistance state for individual switches is given by:
$r_{\text{up}}\propto e^{-(E_{a}-qV/2)/k_{B}T}$
while for the transition from the low resistance state to the high resistance state we have:
$r_{\text{down}}\propto e^{-(E_{a}+qV/2)/k_{B}T}$
In turn, we can see that subtracting the rates gives us the instantaneous, net rate for movement from a high resistance to a low resistance state:
$r_{\text{net}}$ $\propto e^{-(E_{a}-qV/2)/k_{B}T}-e^{-(E_{a}+qV/2)/k_{B}T}$ (5)
$\propto e^{-E_{a}/k_{B}T}\sinh{(-qV/2k_{B}T)}$ (6)
This is similar to the equation for high field ionic drift for the rigid point ion model described in [13]. However, the net rate is only an instantaneous rate of change. In our model, any event (transition from up to down, or down to up) causes a change in the number of switches in each state, and thus a linear and proportional change to the rates $\lambda_{\text{up}}$ and $\lambda_{\text{down}}$. As a result, we must re-calculate the rates whenever there is a change in the state.
II-E Volatility State Variable
Memristive systems, such as PCM and Metal Oxide devices, differ significantly in their structure, material composition, and switching/conduction mechanics, but can nevertheless exhibit similar characteristics. For example, in the case of their switching behaviour under bias, both types of device have associated physical phenomena which cause the device’s state to autonomously change under zero bias, following initial programming. This motivates the introduction of a unified way of describing these disparate physical phenomena for the purposes of modelling.
We introduce the notion of “volatility”. The term “volatility” is sometimes used to refer to the phenomenon of resistive drift - the tendency for the resistance of a device to deviate from its initial value over time. We here generalise the term to include any kind of transient physical phenomenon that can affect the switching or conduction of a memristive device. Examples of physical phenomena that fall under the umbrella of volatility include:
  1. 1.
    Joule heating resulting in nonlinear switching behaviour and/or accelerated carrier drift under bias in various memristive device types, including metal oxide and PCM devices.
  2. 2.
    Structural relaxation in PCM devices.
  3. 3.
    Structural disruption that causes reconfiguration in diffusive memristive devices.
The presence of volatile mechanics in a device tends to express itself in several ways: memory-dependence in switching and conduction behaviour; self-excitatory or self-inhibitory behaviour during switching and conduction leading to non-linear switching and/or conduction; and autonomous state evolution as a result of temporary changes within the devices, such as increased temperature or disorder in the material. The unification of the phenomena removes the need to identify the exact cause of memory-dependent switching behaviour in memristive devices, allowing us to focus only on the consequences for the evolution of the state variable. The volatility state variable summarises the device history that is relevant to future switching behaviour into a single state variable.
In the general sense of the word, an increase in volatility is associated with an increase in the state transition probabilities (in the discrete setting). This contrasts with a directed bias, which will result in an increase in one or more switching rates at the expense of decreasing one or more other switching rates.
Intuitively, if we imagine switching events to be associated with a change in the state of the device, an increase (decrease) in volatility means that the number of state change events per unit time increases (decreases). Depending on the degree to which the different switching rates increase, relative to one another, a change in the volatility may also have an impact on the local metastable (equilibrium or capturing) state(s).
Note that our definition of volatility motivates the way in which we model its effect mathematically for a given switching probability model. For example, in later sections we will explore the use of a Boltzmann activation energy model, and the volatility will have a similar effect to the temperature, increasing all switching rates simultaneously.
II-E1 Phenomena Captured by the Volatility State Variable
One key observation is that volatility is higher immediately after a programming event in a range of memristor devices, including ECM and VCM ionic movement memristors, and in PCM devices. It may also be the case that the degree of volatility is dependent in some way on the nature of the programming (i.e., it is not a memoryless phenomenon given the memristive state alone), with more extensive programming pulse application or a higher deviation from a previous stable level resulting in a higher degree of volatility. Volatility then tends to relax to a stable level in the absence of any stimulating voltage input (or change in ambient conditions such as temperature).
In the examples below, we detail experimental evidence that further justifies and supports the theoretical motivation for using a framework involving volatility as a state variable.
Joule Heating
In the case of filamentary switching behaviour in metal oxide memristors, the drift of oxygen vacancies is often non-linear (exponential) with respect to the applied bias. This is attributed to multiple possible phenomenon, including high-magnitude local electric fields, and possible accelearation through Joule heating effects within the device. It is therefore the case that Joule heating may produce memory-dependent switching behaviour, with the internal temperature of the device - itself affected by the history of voltage applied across the device - affecting the rate of switching.
When the device is being programmed, energy is dissipated and held as heat energy in the device. Over time, if there is no stimulating voltage, Newton’s law of cooling applies and the device returns to the ambient temperature. This is a dynamic process, and the cooling may not be fully complete before subsequent programming pulses. Cooling also occurs during the application of programming pulses, with cooling happening more rapidly, the higher the internal temperature of the device.
The actual picture governing volatility may not be fully described through the consideration of Joule heating effects. Wang et al. have examined the properties of temperature and diffusive dynamics for diffusion-driven, volatile tantalum-oxide memristors [14]. They state that temperature cannot be the driving factor in state relaxation for these diffusive, uni-polar devices, since temperature increases produced by Joule heating relax on the order of $1/\kappa$, whereas the memristor relaxation (stochasticity) occurs on a much longer order of $10-201/\kappa$. Thus, they hypothesise, diffusion is the main driving factor, rather than a remnant increase in temperature. This is applicable to the devices studied, which are based on the formation of conductive bridges formed from silver metal nanoparticles through $SiO_{x}N_{y}$ dielectric substrate doped with silver ions, and sandwiched between $Ag$ electrodes. Their conclusion is that Joule heating’s role in the switching process, or subsequent relaxation (switching in the absence of bias) can vary dependent on the inherent volatility of the device.
Structural Relaxation in PCM
Structural relaxation is a phenomenon in PCM (e.g. GST), that results from the gradual consumption of point defects in the amorphous structure over time, following heating followed by rapid cooling that results in a glass-like structure.
The number of unconsumed defects in the device can be thought of as the level of volatility, which is highest after programming when the device has not had time to partially relax back into a crystalline configuration. As the defects are consumed, the volatility decreases. A more detailed description is as follows: after programming, the device relaxes towards the ideal glass configuration, through the consumption of defects in amorphous GeTe that cause the amorphous phase to differ significantly from the crystalline phase. Although drift does not cause the phase to become crystalline, it is generally agreed that defects in the amorphous GeTe are consumed, moving the bond network towards ’structures with chemical order and coordination numbers similar to those of the crystalline phase’. Note that we emphasise that the structure is similar to the crystalline phase, but it is not identical. True relaxation to the crystalline phase would cause full restoration of the high resistive state [15], which is possible if the device is not programmed fully to the amorphous phase using partial SET operations [16].
Electrochemical Metallization: Entropy and Disruption of Structure
In electrochemical metallization (ECM) devices, the entropy of the oxygen vacancies in the device may also prompt a memory-dependent response. Minimisation of interfacial energy may prompt reconfiguration of metal ions within the insulator bulk for diffusive memristors. This phenomenon is described for the case of diffusive memristors in [14].
The general form of volatility is a state variable which directly modifies the global rate of switching, whose increase is a function of external inputs such as voltage, or current, and which has a characteristic decay time constant, $\tau_{x}$. Example implementations of volatility are given in Section IV-D3, where the effects of both Joule heating/Newton cooling, and of arbitrary volatile effects are considered.
II-F Practical Simulations
Let us consider the switching of a single metastable switch to be modelled by a memoryless exponential distribution with the rate parameter $\lambda$ proportional to the probability of switching. Note that this model is memoryless - the distribution for the evolution of the resistive state can be completely characterised by the current state, with no reference to the previous states of the device.
II-F1 Point Process Sampling
Considering the cumulative distribution function for this exponential distribution, we can use use inverse transform sampling (the Smirnov transformation) to generate samples for the point process:
$p(T\leq t)$ $=1-e^{-\lambda t}$
$y$ $=1-e^{-\lambda t}$
$t$ $=-\frac{\ln{(1-y)}}{\lambda}$
We can perform the simulation of parallel point process sampling in two ways. We note that we can sample the time from a point process with rate equal to the sum of all rates of the independent point processes, and subsequently label the switching events post-fact. Alternatively, we can sample times for a cumulative “up“-switching process and a cumulative “down”-switching process and take the smallest inter-arrival time as the next event, and repeatedly resample.
II-F2 Changes to the Distribution at Events
When an event occurs (either a switching up or switching down event), this impacts the distributions and thus we need to reevaluate a sample for the switching time (for both the upwards and downwards processes).
We use the memoryless property of the exponential distribution, which states that if an event has not occurred by time $\tau$ then the distribution for an event to occur at time $t+\tau$ is still exponentially distributed. The only distinction is that we now modify the rate parameter for each distribution, which can be calculated as a simple linear scaling of the rate parameter, since the sum of N independent and identically distributed exponential distributions with rate parameter $\lambda$ is an exponential distribution with rate parameter $N\lambda$, as described in Section II-D.
II-F3 Simulation Algorithm
The procedure for running the model simulation is given in Algorithm 1. $R$ corresponds to the array of generated resistances (with $r_{\text{init}}$ being the initial resistance, determined by the initial state parameter $n$) and $T$ corresponds to the array of times at which the resistance changes. $\sim$ denotes sampling from a given probability distribution. $f_{\text{ro}}$ is the readout equation that gives us a resistance value for a given state, as detailed in Equation 10.
Note that in the case the value of $n$ falls to zero, or reaches $N$, we simply wait for an appropriate event to occur, to increase or decrease the value of $n$, respectively.
Non-Event-Based Inputs
In order to incorporate non-event based inputs, including variables that may be continuous functions of time such as temperature or applied bias, or piecewise constant functions, we can sample these using threshold-based, or event-driven, sampling. For example, the volatility state variable, introduced in Section II-E, is a continuous function of time. Given an independent input signal $m_{i}(t)$, we can define the associated marked point process, defined by the set of times $\mathcal{T}_{i}$, at which the events occur, to be:
$\mathcal{T}_{i}$ :=$:=\{t:f_{i}(t)\in\mathcal{M}_{i}\}$
where, for continuous functions the set $\mathcal{M}_{i}$ contains the function values, changing in steps of $a$:
$\mathcal{M}_{i}$ =$=\{f(0),f(0)+a,\cdots,\lim_{k\rightarrow\infty}f(0)+ka\}$
$\cup\{f(0),f(0)-a,\cdots,\lim_{k\rightarrow\infty}f(0)-ka\}$
In contrast, for piecewise constant (discontinuous) functions, $\mathcal{M}_{i}$ will simply be the set of associated amplitudes at times $\mathcal{T}_{i}$, known ahead of time.
Due to the memoryless nature of the Poisson process, incorporating such sampling into the model is straightforward - whenever $t\in\mathcal{T_{i}}$, for a particular associated set of events $i$, we enforce the event and then resample times to firing for the Poisson switching processes. The canonical example of an input that can be modelled in this way, is the driving input voltage, that takes on predefined values at predetermined points in time.
Although the method of conversion for continuous functions for inclusion in the event-based modelling framework works to minimise the error associated with changes to parameters affecting the rate equations not matching those used during the simulation - which are dependent purely on external stimuli - additional challenges are presented in the case of variables that are instead, or also, functions of the device state themselves. For example, if Joule heating is dependent on the current through the memristor, then this variable must be re-evaluated whenever there is a change in the resistance, or in the voltage. As such, for variables in the rate equations that are functions of both external stimuli and the internal state variable, these should be re-evaluated every time any event happens, so as to avoid errors associated with the drift of the independent variables of the function away from the values used for the calculation.
We can incorporate these variables by considering them to be a subset of the variables by allowing the updates to be governed by function evaluation at pre-defined times, $\mathcal{T}_{i}$, as described previously, in order to account for external (autonomous) changes to the variable, in addition to at any point when a dependent variable of the function changes. Let $\{j_{0}\cdots j_{j}\}\in\{m_{0}\cdots m_{i}\}$ be the set of functions which depend on at least one other variable. We note that the up and down switching rates themselves, $\lambda_{\text{up}}(n)$ and $\lambda_{\text{down}}(n)$, are within the subset $\{j_{0}\cdots j_{j}\}$, since they are updated whenever the variable $n$ is modified.
The algorithm for switching, including the incorporation of additional updates to parameters which are dependent on external stimuli - not limited to state switching events - is given in Algorithm 1.
Algorithm 1 Event-based modelling procedure, incorporating additional non-event-based inputs.
1:$R\leftarrow R_{\text{init}}$
2:$n\leftarrow n(R)$
3:$M_{i}\leftarrow\mathcal{M}_{i}[0]$
4:$\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(\{M_{i}\})$
5:$\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(\{M_{i}\})$
6:$c_{i}\leftarrow 0$
7:$t\leftarrow 0$
8:$F\leftarrow 0$triangleright State variable which denotes whether a firing event has occurred.
9:while $t\leq\Delta T$ do
10:     $t_{\text{down}}\sim\text{Exp}((N-n)\cdot\lambda^{\prime}_{\text{down}})$
11:     $t_{\text{up}}\sim\text{Exp}(n\cdot\lambda^{\prime}_{\text{up}})$
12:     for all $i$ do $\triangleright$ Firing for the state-independent variables
13:         if $F=0\;\textbf{and}\;c_{i}\leq|\mathcal{T}_{i}|\;\textbf{and}\;\mathcal{T}_{i}[c_{i}+1]\leq\{t_{\text{up}},t_{\text{down}},\mathcal{T}_{k}[c_{k}+1]\;\forall k\neq i\}$ then
14:              $c_{i}\leftarrow c_{i}+1$
15:              $t\leftarrow\mathcal{T}_{i}[c_{i}]$
16:              $M_{i}\leftarrow m_{i}({M}_{i}[c_{i}],n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
17:              $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
18:              $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
19:              $F\leftarrow 1$
20:     if $F=0\;\textbf{and}\;(t_{\text{down}}\leq t_{\text{up}}\;\textbf{or}\;n=N)\;\textbf{and}\;n\neq 0$ then
21:         $n\leftarrow n-1$
22:         $t\leftarrow t+t_{\text{down}}$
23:         for all $j$ do $\triangleright$ Update the state-dependent variables
24:              $J_{i}\leftarrow j_{i}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
25:              $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
26:              $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
27:     else if $F=0\;\textbf{and}\;(t_{\text{up}}<t_{\text{down}}\;\textbf{or}\;n=0)$ then
28:         $n\leftarrow n+1$
29:         $t\leftarrow t+t_{\text{up}}$
30:         for all $j$ do
31:              $J_{i}\leftarrow j_{i}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
32:              $\lambda^{\prime}_{\text{up}}\leftarrow f_{\text{up}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
33:              $\lambda^{\prime}_{\text{down}}\leftarrow f_{\text{down}}(n,M_{0},\cdots,M_{i},J_{0},\cdots,J_{j})$
34:     $R\leftarrow R(n)$
35:     $F\leftarrow 0$
Conversion to Discrete Time
If necessary, we can convert non-uniform series from the event-based model to uniform data in time through sampling, albeit introducing errors associated with the discretisation in time, as shown in Algorithm 2. $\mathcal{T}_{switch}$ denotes the ordered set of switching times, and $\mathcal{R}$ denotes the associated ordered set of values of a state variable at the times $\mathcal{T}_{switch}$. $T_{\text{sample}}$ denotes the sampling period and $T_{\text{tot}}$ denotes the maximum desired time to sample for.
Algorithm 2 Procedure for conversion of the non-uniform time series to uniformly sampled time series.
$i\leftarrow-1$
2:$t_{\text{curr}}\leftarrow 0$
3:$\mathcal{R}_{\text{sample}}\leftarrow\{\}$
4:for all $t\in\mathcal{T}_{switch}$ do
5:     if $t\geq T_{\text{curr}}$ then
6:         for all $n\in[1,1+\lfloor\frac{t-t_{\text{curr}}}{T_{\text{sample}}}\rfloor]$ do
7:              $\mathcal{R}_{\text{sample}}\leftarrow\mathcal{R}_{\text{sample}}\cup\{\mathcal{R}[\max{(0,i)}]\}$
8:              $t_{\text{curr}}\leftarrow t_{\text{curr}}+T_{\text{sample}}$
9:         $i\leftarrow i+1$
10:         if $t_{\text{curr}}>T_{\text{tot}}$ then break
11:     else
12:         $i\leftarrow i+1$
13:if $\lfloor\frac{T_{\text{tot}}-t_{\text{curr}}}{T_{\text{sample}}}\rfloor>0$ then
14:     for all $n\in[1,\lfloor\frac{T_{\text{tot}}-t_{\text{curr}}}{T_{\text{sample}}}\rfloor]$ do
15:         $\mathcal{R}_{\text{sample}}\leftarrow\mathcal{R}_{\text{sample}}\cup\{\mathcal{R}[\max{(0,i)}]\}$
16:$\textbf{return}\;\mathcal{R}_{\text{sample}}$
III Model Analysis
III-A Model Analysis
Our model has thus far been defined, but we would also like a computationally efficient way of evaluating it. Our model is a marked [^2] A point process whose events are associated with additional information, such as amplitude. In this case - the direction of switching of the state variable. Hawkes [^3] A point process whose events impact the rate parameter (traditionally self-excitatory, but also allowing for self-inhibitory behaviour). process with both self-excitatory and self-inhibitory behaviour. We would like to demonstrate different ways of looking at the behaviour of the model, in order to motivate the development of efficient schemes for numerical simulations and prediction of the statistics.
The majority of our analysis in this section will be under the assumption of static environmental conditions (temperature and bias), since under varying conditions, additional modifications to the switching parameters, and thus, to the equilibrium point of the model, will occur. The change of the memristive state in the absence of a voltage bias is usually referred to as the phenomenon of resistive drift, or volatility. The equilibrium point of the model under steady state conditions is identified in Section III-A1.
III-A1 Calculation of the Equilibrium Point (Metastable Point)
Given the forward switching rate, we can calculate the equilibrium number of switches in the low resistance state at a given time by equating $\lambda_{\text{down}}$ to $\lambda_{\text{up}}$. The expression for this is given in Equation 7, below. A point of dynamic equilibrium is attained when the rate $\lambda_{\text{down}}=\lambda_{\text{up}}$; that is, the rate of switching from the high resistance state to the low resistance state is equal to the rate of switching in the opposite direction. In the case that we have a number of states greater than 2, the determination of the equilibrium state is less straightforward and the steady state vector of the associated Markov state transition matrix must be identified.
$\lambda_{\text{up}}$ $=\lambda_{\text{down}}$
$n_{\text{eq}}e^{-\frac{E_{a}-V(t)q/2-E_{\text{off}}/2}{k_{b}T(t)(1+\rho(t))}}$ $=m_{\text{eq}}e^{-\frac{E_{a}+V(t)q/2+E_{\text{off}}/2}{k_{b}T(t)(1+\rho(t))}}$
$n_{\text{eq}}e^{\frac{2(V(t)q/2+E_{\text{off}}/2)}{k_{b}T(t)(1+\rho(t))}}$ $=m_{\text{eq}}$
$n_{\text{eq}}$ $=\frac{N}{e^{\frac{(V(t)q+E_{\text{off}})}{k_{b}T(t)(1+\rho(t))}}+1}$ (7)
We can see that the equilibrium point, as a fraction of $N$, is independent of the value of $E_{a}$, depending only on $E_{\text{off}}$ (as well as $T(t)$, $V(t)$, and $\rho(t)$). This shows that the equilibrium point and the rate of attainment of the equilibrium point can be independently adjusted, using $E_{\text{off}}$ and $E_{a}$, respectively.
The equilibrium resistance can be determined from the overall equilibrium state via the readout equation.
When external environmental parameters are functions of time, the model’s equilibrium point and the rate parameters associated with the switching point processes will also be functions of time, changing even when there are no underlying state changes due to metastable switching events.
Under a zero- or a piecewise constant- bias setting, the phenomenon known as resistive drift occurs for memristors. This can be shown to be a competition between two Poisson processes with linearly varying rates. There is a well-defined equilibrium point towards which resistive drift occurs, which can be calculated by finding the state for which the rates of the competing switching processes are equal. In the steady state, in a constant-bias setting, the temperature will no longer be time varying. In a zero-bias setting specifically, the temperature will be equal to the ambient temperature. The zero-bias equilibrium point therefore has a simplified expression given as follows:
$n_{\text{eq}}$ $=\frac{N}{e^{\frac{qV_{\text{off}}}{k_{B}T}}+1}$ (8)
We can express the offset voltage in terms of $n_{\text{eq}}$ as:
$V_{\text{off}}$ $=\ln{\left(\frac{N-n_{\text{eq}}}{n_{\text{eq}}}\right)}\cdot\frac{k_{B}T}{q}\equiv\ln{\left(\frac{m_{\text{eq}}}{n_{\text{eq}}}\right)}\cdot V_{T}$ (9)
where $V_{T}=\frac{k_{B}T}{q}$ is a constant, often referred to as the thermal voltage.
III-A2 Additional Discussion
Impact of the Paramter $N$
One strength of a discrete modelling strategy, with a finite number of possible system states (in our case, $N$), is that simulations can be conducted at different resolutions, depending on precision requirements, and the scale of interest. For example, if a memristive device is to be used as a binary storage device, where only the states $1$ and 00 are distinct, then a value of $N=2$ suffices to model the readout value of interest. In contrast, for a simulation where multiple stable states are to be for information storage, a larger value of $N$ may be required. The “best” value of $N$ to use is a matter of perspective - it depends on the extent to which regions of device resistance or behaviour are considered to be equivalent for the intended purpose, as exemplified by the binary storage example, where the many possible configurations of the device are summarised by a single bit (2 possible values) of information.
As a consequence of the fact that the variance and mean of a Poisson process are equal to the rate of the process, the magnitude of the parameter $N$ will scale the expected value and the variance of the final state in a linear fashion. As a result, the standard deviation in the final state, after a given delay, is proportional to the square root of $N$. Thus, we can achieve less stochastic switching by scaling up $N$. Achieving the same mean, for a given delay, for the readout variable can be achieved by applying a straightforward linear scaling of the parameter $n$ in the readout equation. When we increase N, the average jump time decreases linearly with N (the rates of up and down jumps increase proportional to N). This means that over the timescale of interest, many more trials occur, with each trial also contributing to a smaller percentage change in the resistance, resulting in a decrease in the variance with increasing $N$. Continuous and deterministic switching behaviour can be approximated in the limit as $N\rightarrow\infty$.
Error Compared to Timestep-Based Method
Note that the error, in contrast to the GMSM, is related to the level of quantisation and the type of quantisation. In the case of the previous method, we have errors related to missing switching events, which impact the instantaneous probability distribution. We also have error associated with the approximation of a sum of binomials by a normal distribution. The normal distribution’s parameters at each point in time will change with each switching event; however, this is not reflected by the choice of the use of a constant delay between each distribution update. The duration of the delay chosen will have an impact on the accuracy as well.
In contrast, in the case of event-based switching, no error is introduced if the parameter N accurately reflects the physical mechanism. Using a value of N that is smaller will introduce a quantisation error (but no bias) to the switching process evaluation. That is - the expected value of the process after given number of switching events will be the same, since the value of $n_{eq}$ scales linearly with $N$, and the associated parameters of the readout equation can be adjusted accordingly.
Note that integrate and fire sampling for continuous functions of time, as described in Section II-F3, will unavoidably introduce discretisation errors. If the functions vary slowly in time, then the number of additional events due to the introduction of additional functions will be small. However, if the functions vary quickly in time and the number of generated events is large, then the benefits of event-based modelling approaches diminish, and the model’s events become dense in time. This still, however, offers benefits over a uniform sampling approach, since it associates the density of samples in time with the rate of information change that impact the model updates (and hence the error associated with discretisation in time).
IV RRAM Metal Oxide Device Example
In this section, we look at an example of a titanium dioxide memristive device to model the resistive drift phenomenon under zero bias.
IV-A Physical Model
Modelling of metal oxide metal-insulator-metal (MIM), and more specifically - Titanium Dioxide MIM, devices has been studied previously. Titanium Dioxide memristor switching has been previously modelled as a single conductive pathway of varying resistance [5], and other similar devices, such as tantalum oxide memristors have been modelled as exhibiting the dominance of a single filament or conductive region [17].
Based on initial investigation of a drift dataset (see Section IV-C below), we model the conductance of the device as linearly proportional to state variable $n$. Based on this information, we hypothesise that the mechanism underlying the switching of the device modulates the width of a conductive region within the active substrate, bridging a connection betwen the electrodes. Phenomena that could account for this behaviour include a single conductive filament that grows or reduces in width, or multiple conductive filaments that each contribute linearly to the conductance of the device. We observe that it is necessary to use a value of $n_{\text{extra}}$ and $m_{\text{extra}}$ much greater than the value of $N$. This suggests a thresholding effect, where the rate of switching is relatively constant in the switching region, under a constant voltage bias. This corresponds somewhat to the ideal behaviour for a memristor, which is why such titanium dioxide devices were presented as candidate memristive devices originally [5]. We assume a large $n_{\text{extra}}=10\times N$ such that the switching rates remain relatively constant across the considered range of resistance values.
Our model suggests that a large number of switches must be in the low resistance state before there is any measurable impact on the rate. Switching a sufficient number of the switches into the low resistance state can be thought to be performed during the electroforming step. We can imagine our model being a conductive filament that grows in length, eventually making physical contact with the electrode on the other side of the active layer, allowing electrical conduction to occur. Once this contact occurs, additional switching causes the contact region to grow linearly in area of contact, resulting in a linear increase in the conductive state once we are in this “active” region. The model for this process is visualised in Figure 4.
Figure 4: Our physical model for switching in a titanium dioxide memristor. The growth of the filament will not significantly impact the resistance (conductance) - i.e. the readout value - until the filament size crosses a particular threshold. Once this condition is achieved (middle figure), we are in the active region of switching, where small changes in the state variable have a significant impact on the resistance, due to the modulation of an effective area of conduction through the active layer of the device.
IV-B Readout Equation
In the case of Titanium Dioxide devices, we assume that the state variable constitutes the number of parallel filaments that have been formed through the device, or alternatively, the diameter of a single conductive filament. In both cases, the resistance will not be linearly related to this state variable. Rather, it is the conductance, thus our readout equation will be:
$R(n)=\frac{1}{g_{\text{diff}}\max{(n,n_{\text{thresh}})}+\frac{1}{R_{\text{high}}}}$ (10)
We can see from the equation that the behaviour of the model is thresholded, with the resistance remaining at the value $R_{\text{high}}$ until the value of $n$ surpasses the parameter $0\leq n_{\text{thresh}}\leq N$, upon which the conductance increases linearly in steps of $g_{\text{diff}}$. This is motivated by the conductive filament model described above.
We can also invert the equation to find $R$ in terms of $n$:
$n(R)=n_{\text{thresh}}+\frac{1}{g_{\text{step}}}\left(\frac{1}{R}-\frac{1}{g_{\text{parallel}}}\right)$ (11)
This equation is useful for determining the correspondence between a certain resistance value and the number of switches in the “low” state, $n$, for example, for determining the desired $n$ value for a specified equilibrium resistance. Note that if the resistance is at its maximum, we simply assume the value of $n$ to be $n_{\text{thresh}}$, although it could be any value at or lower than this. This is based on the assumption that we lie in or close to the active switching region once the device has been electroformed.
Note that though the readout equation is not linear in the state variable, the equivalent conductance (reciprocal of the resistance) is linear in the state variable, the physical motivation for which is provided above. From the readout equation, we can see that the active region will begin at approximately $n=n_{\text{thresh}}$. This is visualised in Figure 5.
Figure 5: The active switching region is defined and shown as the region where changes to the state variable measurably impact the readout variable (in this case the resistance). Before this point, changes to the state are modelled not to have an impact.
IV-C Dataset
To evaluate the behaviour of the model under controlled conditions, we created a dataset based on the evolution of memristive states. We collected 2751 device hours of data across two separate memristor arrays. We set each of the devices in the array to a random initial resistance and measured the evolution of the resistance over time, using a short 30ms pulse of 0.3V magnitude to read the instantaneous resistance every 10 seconds. In total, 208 data series were collected, of varying length. The devices used were thin-film Titanium Dioxide ($TiO_{2}$) memristors with $Al_{2}O_{3}$ interstitials (the same as [5]), placed on the diagonal of crossbar arrays to mitigate sneak path currents. The measurements were collected using the ArCONE memristor characterisation platform [18].
After collection, the series were inspected and a data cleaning step was performed. As part of this step, series for memristors which had failed or degenerated to a non-volatile state after programming, as well as clearly anomalous measurements within series due to measurement errors, were removed. Following data cleaning, 14 series were excluded from the dataset, leaving 196 series.
IV-C1 Quantisation
Figure 6 shows a histogram of the unquantised resistance values in the dataset, showing the range of interest to be the interval of approximately $[10^{3},10^{5}]$, where changes to the value of $n$ significantly affect the output resistance. As such, we choose the values of $N$, $n_{\text{thresh}}$, $g_{\text{step}}$, and $g_{\text{parallel}}$ accordingly. We set $g_{\text{step}}=10^{-7}$, $g_{\text{parallel}}=10^{-10}$, and $N=2n_{\text{thresh}}=20,000$.
Figure 6: The unquantised resistance dataset, showing the empirical distribution of all resistance values in all series.
After choosing the parameter values that determine the number of resistance states and their values, we calculate the resistance corresponding to each possible value of $n$, using the readout equation given in Equation 10. Let $R(n)$ be the readout equation mapping the state variable $n$ to the corresponding resistance value. The resistance bin boundaries, $\mathcal{R}_{\text{bound}}$ and values $\mathcal{R}_{\text{val}}$, in the order of decreasing resistance (increasing state), are given by:
$\mathcal{R}_{\text{bound}}=$$\{-\infty,R(n_{\text{thresh}}+0.5),$
$R(n_{\text{thresh}}+1.5),\cdots,R(N-0.5),\infty\}$
$\mathcal{R}_{\text{val}}=$$\{R(n_{\text{thresh}}),R(n_{\text{thresh}}+1),$
$\cdots,R(N-1),R(N)\}$
where $\begin{cases}|\mathcal{R}_{\text{bound}}|&=N-n_{\text{thresh}}+2\\ |\mathcal{R}_{\text{val}}|&=N-n_{\text{thresh}}+1\end{cases}$
IV-C2 Pairs
For each series of quantised resistance values in the dataset, we obtain pairs of resistance values from the sequential, non-overlapping intervals of length $T=10,000$ seconds (2.778 hours). This yields a dataset consisting of pairs of initial and final resistance values for the given interval, allowing us to evaluate the statistics of the resistance drift for such an interval, conditioned on the initial resistance value.
IV-D Model Fitting
IV-D1 Linear Conductance Approximation
Based on the physical model described above, we use the approximation that the conductance changes linearly, at a rate proportional to the time. We verify this by measuring the mean change in the conductance conditioned on each of the initial quantised resistance values, $\mathcal{R}_{\text{val}}$, we calculate the mean change in the resistance across all the pairs in the dataset for given initial resistances. A plot showing this is shown in Figure 7.
Figure 7: The mean change in resistance for the dataset of pairs with $\Delta T=10000$. We show a moving average of the data with different window sizes (5 and 50), for a clearer view of the non-linear trend.
We use a zero-bias model consisting of an approximately constant change in the conductance over the given time interval, that is independent of the initial resistance. This assumption is based on our own model, which fits a trend in the observed data to a reasonable standard. We use an iterative trust region fitting algorithm [19] to minimise the squared error associated with fitting the function to the predicted resistance change. The model is given as follows, with parameter $a$ the change in the state variable - fit to the data:
$\Delta R(R_{\text{init}})=\frac{1}{g_{\text{parallel}}-ag_{\text{diff}}}-R_{\text{init}}$ (12)
For the chosen timestep of $10,000$ seconds, we find a value of $a=3.62$ yields the best fit, for the given parameters $g_{\text{diff}}$, $g_{\text{parallel}}$, $N$, and $n_{\text{thresh}}$. This associated line of best fit for this linear conductance change model is shown in Figure 7. The moving averages are plotted to enable a clearer view of the general trend. Although their shape is clear, they do include bias associated with the (symmetric) averaging window size, with a larger window size tending to overestimate the mean change, as can be seen from the larger magntiude of the moving average plot for a window size of 50, compared to the model fit to the data. The moving average window of size 5 yields a (stochastic) function whose mean closely matches the model.
IV-D2 Parameter Fitting and Constant Rate Approximation
We would like to obtain parameters that allow the mean behaviour of the model to correspond to the mean predictions of the linear conductance model. Due to the difficulty in obtaining an analytic expression for the delay-conditional distribution, we use a constant rate approximation to obtain approximate mean values for this distribution over time. We assume that the up and down switching rates of the model remain relatively constant. In order to fit the model, we calculate the rate at a value of $n=n_{\text{thresh}}$ and make the assumption that the rate remains equal to this rate over the region of interest. Thus, this approximation is only valid if the quantities $\frac{n_{\text{thresh}}-a}{n_{\text{thresh}}}$ and $\frac{N-n_{\text{thresh}}+a}{N-n_{\text{thresh}}}$, approximating the percentage change in the up and down switching rates remain relatively close to 1.
We choose the parameters $N$ and $n_{\text{thresh}}$ accordingly. As such, we require $n_{\text{thresh}}>>a$ and $N-n_{\text{thresh}}>>a$. We must also ensure that $N+1-n_{\text{thresh}}$, defining the number of unique resistance states according to Equation 10, is appropriately selected to give a sufficiently fine quantisation over the switching region for which we have data.
For the given value of the parameter $a$, we can determine an appropriate value of $V_{a}$. We first select a value of $V_{\text{off}}$ to ensure an equilibrium point (see Section III-A1) that is far above the active switching region (in terms of resistance), to ensure that the assumption that the rate is constant over the considered range holds, since the equilibrium point at a particular bias sets the zero-crossing point for the rate equation, hence our constant rate assumption will break down as the value of $n$ approaches $n_{\text{eq}}$. In our case, this corresponds to setting an equilibrium point with a value of $n<<n_{\text{thresh}}$. We can see in Figure 8 how the choice of $V_{\text{off}}$ impacts $n_{\text{eq}}$; and hence, the constant rate assumption, and the fit to the linear conductance model. Assuming a constant rate and a given value of $V_{\text{off}}$, we can derive a value of $V_{a}$ according to the following equation:
$V_{a}=-\frac{k_{B}T}{q}\ln{\left(2\Delta Tn_{\text{thresh}}\sinh{\left(V_{\text{off}}\frac{q}{2k_{B}T}\right)}\right)}$ (13)
Figure 8: Fitting the event-based model to the linear conductance model. We use a constant rate assumption to derive approximate values for the parameters.
For very small values of $V_{\text{off}}$, the constant rate assumption will not hold, since the equilibrium point of the model will be high and closer to the active region.
In Figure 9, we see averaged zero-bias simulations from the event-based model, based on the fit parameters, showing the mean change in the resistance fitting to the linear conductance model, as well as the associated variance of the evolutions (which increases with the magnitude of the resistance).
Figure 9: The mean change in the resistance following fitting to the linear conductance model, with errorbars indicating the standard deviation of series, for simulations with length $\Delta T=10000$.
IV-D3 Volatility
In the case of thin film memristors, volatility can be decomposed into two variables. The first of these is the temperature (that may increase due to Joule heating) and the second is a dimensionless volatility state variable to account for additional volatile effects, such as disruptions to the internal structure of the memristor on application of a bias voltage and subsequent restoration of order following its removal (interfacial entropy minimisation).
In modelling the effects of Joule heating, we can assume that a proportion of the current flowing through the device (proportional to the square of the current flowing through the device) is converted to heat energy and raises the internal temperature. We can also assume that Newton’s law of cooling applies simultaneously, resulting in power dissipation and restoration of the internal temperature to the equilibrium (bath) temperature.
To introduce Joule heating as a form of volatility, one can define the rate of change of temperature in terms of a thermal time constant that is a function of a thermal resistance ($R_{th}$) and thermal capacitance ($C_{th}$) [20], dictating how easily heat energy in the form of temperature captures and is dissipated from the device, wit the thermal time constant denoted $\tau_{th}=C_{th}R_{th}$. The differential equation for the internal temperature of the device is given as:
$\frac{dT}{dt}=\left(T_{\text{bath}}+R_{th}\frac{V^{2}(t)}{R(n)}-T\right)/\tau_{th}$ (14)
where $R(n)$ is the instantaneous resistance of the memristor (as a function of the state variable, $n$).
As an example of a form of additional volatility due to disruption of the structure, we assume that the voltage applied results in a bias-dependent equilibrium volatility, and the volatility approaches this value according to a differential equation, where the rate of change of the volatility is dependent on the difference between the instantaneous volatility and the equilibrium value. The volatility and its evolution is defined by the following differential equation:
$\frac{d\rho(t)}{dt}=\left(c_{\text{volatile}}\cdot V(t)-\rho(t)\right)/\tau_{\text{volatile}}$ (15)
where $\tau_{\text{volatile}}$ is the volatility time constant, dictating how quickly the equilibrium value is attained (and therefore how slowly a non-zero volatility decays in the absence of an input signal); and $c_{\text{volatile}}$ is the volatility factor, which - when multiplied by the applied bias - yields the equilibrium volatility (dictating the volatility state variable’s leverage on the switching rates).
IV-E Simulation
In addition to demonstrating the fit of the zero-bias mean values to those determined according to the linear conductance model, we demonstrate the capabilities of the model to simulate memristive behaviour through a series of simulated experiments for a titanium dioxide thin film memristive device.
IV-E1 Experimental Parameters
The parameters used for the experiments, chosen for $\Delta T=10000$, are summarised in Table I. The choices for these parameters are given in Sections IV-C1 and IV-D2.
TABLE I: Table of parameters for the model, for titanium dioxide memristive devices.
Parameter Value
$T_{\text{bath}}$ 300
$V_{a}=E_{a}/q$ 0.40049
$V_{\text{off}}=E_{\text{off}}/q$ 0.05
$N$ 20,000
$g_{\text{parallel}}$ $10^{-10}$
$g_{\text{step}}$ $10^{-10}$
$n_{\text{thresh}}$ 10,000
$\tau_{th}=C_{th}\cdot R_{th}$ $3.84\times 10^{-14}\cdot 4\times 10^{4}$
$c_{\text{volatile}}$ 10
$\tau_{\text{volatile}}$ 10
IV-E2 Switching Experiments
We use the model, with the derived parameters, to generate example device data for simulated input signals. Our first experiments assume there is no volatile behaviour. In the second set of experiments, we assume Joule heating - modulated by the current flowing through the device - and associated restoration of the equilibrium temperature according to Newton’s law of cooling, as described in Section IV-D3.
We perform three sets of experiments: pulsing with positive voltage pulses (increasing/decreasing in magnitude) to increase the device resistance, and with negative voltage pulses (increasing/decreasing in magnitude) to decrease the device resistance. We also examine the effect of alternating pulses. The results of these experiments are shown in Figure 10.
The input signal parameters are given in Table II. We simulate series from starting resistances ranging between $10k\Omega$ and $100k\Omega$. We choose the parameters for the input signals in order to highlight model behaviour. In all the experiments, we see that the model is biased towards movement to a higher resistance state, with negative pulses having a lesser impact on reducing the resistive state than positive pulses have on increasing the state. During the inactive period of each pulse, we see the device’s resistance drift in the upwards direction as a result of the impact of $V_{\text{off}}$.
TABLE II: Parameters for the input signals in the simulated switching experiments. In the case of the positive descending experiment, we observe that too large a signal amplitude drives the memristor immediately to a high resistance state, so we lower the initial amplitude to better observe the behaviour. In the case of the alternating experiment, we use a small duty cycle such that the active pulse duration is 1s.
Experiment Max/Min Pulse Pulse Duty
Amplitude Step Duration (s) Cycle
Negative Ascending -0.2 -0.01 1000 0.1
Negative Descending -0.2 -0.01 1000 0.1
Positive Ascending 0.2 0.01 1000 0.1
Positive Descending 0.07 0.01 1000 0.1
Alternating 0.5 N/A 1000 0.001
Negative Ascending Negative Descending Alternating Positive Ascending Positive Descending
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Simulated experiments for a variety of different stimulating voltage inputs. From the alternating pulses experiment, it is clear that the preferred direction of movement (the equilibrium resistance) being a high value, results in the final resistance being higher than the initial resistance, due to the impact of resistive drift, despite the same number and magnitude of positive and negative pulses being applied. Green lines indicate voltage inputs and red/blue lines are resistance evolutions.
V Conclusions and future work
We have introduced an event-based modelling strategy for memristive devices by extending previous generic modelling approaches for memristors. We introduced an event-based modelling strategy based on Poisson processes which addressed errors associated with discrete time approximation in existing models, as well as being natively suited to neuromorphic applications. We proposed a volatility state variable, to allow for the capture of memory-dependent switching behaviour, motivating this from a physical perspective for various memristive device types. We have also presented a resistive drift dataset, which we use to obtain parameters for an example of the model applied to a real device, along with a new linear conductance model of drift in the devices. We demonstrated the ability of the event-based model to replicate the behaviour under zero bias. Future work may include extending the modelling approaches described to additional real data from different memristive device types, under a variety of additional conditions, as well as exploration of the applications of the model for event-based neuromorphic systems and simulations, especially spiking neural networks implemented in hardware.